# Generative-Enhanced Heterogeneous Graph Contrastive Learning

### v2

@@ 5.5. Interpretability and Visualization (RQ4)
Visualization. To more intuitively perceive the embeddings of different models, we use t-SNE (Van der Maaten and Hinton, 2008) to reduce the dimensionality of the embeddings and visualize them with labels. As shown in Figure 6, we select three representative models: unsupervised MP2vec (Dong et al., 2017), contrastive HeCo (Wang et al., 2021), and generative HGMAE (Tian et al., 2023). We can observe that MP2vec performs the worst, with different types of nodes mixed together and difficult to distinguish. HeCo can effectively differentiate categories, but each block is separated and does not form obvious clusters. HGMAE uses a hidden dimension of 256 and is a generative method, thus exhibiting more evident clustering phenomena. However, our GC-HGNN is based on a contrastive framework and demonstrates both clustering and uniform distribution, balancing the advantages of contrastive and generative approaches. This phenomenon also indicates that our generative-contrastive paradigm is effective.
5.6. Parameter Sensitivity (RQ5)
@@
In this study, we explored the impact of key hyperparameters on heterogeneous graph node classification. Moreover, we investigated the effects of hierarchical contrast coefficient, embedding dimension and mask ratio.
@@
Hierarchical contrast coefficient. As shown in Section 4.4, the $\mathcal{L}_{hcl}$ contains $\lambda_{{Intra}}$ and $\lambda_{{Inter}}$, which are set from 0 to 1 with an increment. In Figure 7, we can observe that except for the Aminer dataset, the performance of other datasets improves as $\lambda_{{Inter}}$ increases. This phenomenon follows the principle of our proposed sampling method, where inter-contrast has more hard negative samples compared to intra-contrast.
@@
Embedding dimension. We investigated how different embedding sizes affect node classification performance from 32 to 512. Figure 8 shows that most datasets improved up to a 512-dimensional embedding, apart from DBLP, which peaked at 256. However, to balance the marginal gains against higher computational costs and overfitting risks, we standardized the embedding size to 64 for all tests.
@@
Mask ratio. In Section 4.2, we use a masked autoencoder, where the mask ratio plays a crucial role as a hyper-parameter in generating negative samples. As shown in Figure 8, our model typically achieves better results with a lower mask ratio, contrary to the conclusions drawn by HGMAE (Tian et al., 2023). We explain that, compared to generative feature reconstruction, the game between contrastive learning and generation learns to reconstruct the original data distribution by minimizing the distributional divergence rather than reconstructing individual samples (Liu et al., 2021).
@@
6. Conclusion
In this paper, we propose a novel generative-contrastive heterogeneous graph neural network named GC-HGNN. It is a generative learning enhanced contrastive paradigm and a meaningful exploration. We combine meta-paths and masked autoencoders to create augmented contrastive views without disrupting the data structure. Then, we propose position-aware and semantic-aware sampling strategies to alleviate the synthesize hard negatives problem in heterogeneous graphs. Finally, we propose a hierarchical contrastive learning self-supervised framework to capture local and global information. In particular, the hierarchical contrastive learning and sampling strategies constitute the enhanced contrastive discriminator under this paradigm. Experiments demonstrate that this contrastive discriminator effectively improves the stability of output embeddings. The generative-contrastive framework significantly improves the performance of node classification and link prediction tasks. We compare our model with seventeen state-of-the-art models on eight real-world datasets. Our model outperforms the latest contrastive and generative baselines on multiple tasks.

### v3

@@ IV-F2 Visualization
To gain a more intuitive understanding of the embeddings from different models, we employ t-SNE [51] to reduce the dimensionality of the embeddings and visualize them with labels. As illustrated in Figure 6, we select three representative models: unsupervised MP2vec [3], contrastive HeCo [7], and generative HGMAE [23]. We can observe that MP2vec performs the worst, with different types of nodes mixed together, making them difficult to distinguish. HeCo effectively differentiate categories, but each block remains separated and does not form distinct clusters. HGMAE uses a hidden dimension of 256 and is a generative method, thus exhibiting more evident clustering phenomena. However, our GC-HGNN, based on a contrastive framework, demonstrates both clustering and uniform distribution, effectively balancing the advantages of contrastive and generative approaches. This phenomenon further indicates the effectiveness of our GC-based paradigm.
IV-G Parameter Sensitivity (RQ4)
@@
In this study, we explore the impact of key hyperparameters on heterogeneous graph node classification. Moreover, we investigate the effects of hierarchical contrast coefficient, embedding dimension, and mask ratio.
@@
  • Hierarchical contrast coefficient. The $\mathcal{L}_{hcl}$ contains $\lambda_{{Intra}}$ and $\lambda_{{Inter}}$, which are set from 0 to 1 with an incremental step. In Figure 7, it is observed that the performance improves as $\lambda_{{Inter}}$ increases. This phenomenon aligns with the principle of our proposed sampling method, where inter-contrast has more hard negative samples.
  • Generative coefficient. In Table IV-F1, we evaluated $\mathcal{L}_{Gen}$ ranging from 0.1 to 0.5, observing that performance metrics typically decline as the $\mathcal{L}_{Gen}$ increases. It can be observed that the model tends to favor a smaller (generative learning) GL loss, which aligns with our idea regarding the auxiliary role of GL in the (generative-contrastive) GC paradigm.
@@
  • Embedding dimension. We examine how different embedding sizes affect node classification performance ranging from 32 to 512. Figure 8 shows that most datasets show improvement up to a 512-dimensional embedding, apart from DBLP, which peaks at 256. However, to balance the marginal gains against higher computational costs and overfitting risks, we standardize the embedding size at 64 for all tests.
@@
  • Mask ratio. We utilize a masked autoencoder, where the mask ratio plays a crucial role as a hyper-parameter in generating negative samples. As shown in Figure 8, our model achieves better results with a lower mask ratio, contrary to the conclusions drawn by HGMAE [23]. We attribute this to the fact that, compared to generative feature reconstruction, the game between contrastive learning and generation learns to reconstruct the original data distribution by minimizing the distributional divergence rather than reconstructing individual samples [14].
  • Encoder types in MAE. We set multiple variants for the encoder and decoder to test their effect on GC-HGNN. Theoretically, our encoder can be any GNN model (such as GIN [52]), or even an MLP. As shown in Table IV-F1, the best result is highlighted in bold, while the second-best is underlined. We choose a combination of the commonly used GAT [29] and GCN [53]. It can be observed that the GAT-GAT combination achieves the best or second-best performance on most datasets. This phenomenon aligns with the conclusion that the attention mechanism can effectively capture relationships between nodes in node classification tasks. Moreover, all variants outperform the MAE-free variant from the ablation study, demonstrating that MAE effectively mitigates noise limitation.
@@
V Conclusion and Future Work
In this paper, we proposed a novel generative-contrastive heterogeneous graph neural network termed GC-HGNN. It represented a generative learning enhanced contrastive paradigm and a meaningful exploration. We combined meta-paths and masked autoencoders to create augmented contrastive views without disrupting the data structure. The generative-contrastive framework significantly improves the performance of node classification and link prediction tasks. We compared our model with seventeen state-of-the-art models on eight real-world datasets, demonstrating that it outperformed the them on these tasks.
Acknowledgment
This work is supported by the National Science Foundation of China (No. 62206002, No. 62272001 and No. 62206004), Hefei Key Common Technology Project (NO. 2023SGJ014), Natural Science Foundation of Anhui Province (No. 2208085QF195) and Xunfei Zhiyuan Digital Transformation Innovation Research Special for Universities (NO. 2023ZY001).